OpenStudy (anonymous):

Write an equation of the sine function with amplitude = 3 and period = 2pi.

OpenStudy (anonymous):

sine oscillates between 1 and 0; its amplitude is 1 to get an amplitude of 3, the sine must be multiplied with A=3

hartnn (hartnn):

general equation of sine function is \(f(x)=A\sin (2\pi x/T+\phi)\) where A = amplitude T = period \(\phi\) - phase shift

hartnn (hartnn):

here, A and T are given, just plug in values! assume phase shift =0

OpenStudy (anonymous):

thanks guys

hartnn (hartnn):

welcome ^_^